A solution containing 30% insecticide is to be mixed with a solution containing 50% insecticide to make 200 L of a solution containing 42% insecticide. How much of the 50% insecticide solution should be used?
Answer by checkley77(12844)

.50X+.30(200-X)=.42*200
.50X+60-.30X=84
.20X=84-60
X=24/.20
X=120 L OF 50% INSECTICIDE IS USED.
PROOF:
.50*120+.30(200-120)=.42*200
60+.3*80=84
60+24=84
84=84